OpenStudy (anonymous):

True or False A. A dead duck is placed on a scale; the scale reads 15 N. A giant vacuum cleaner sucks air from the room, reducing the pressure to 0.8 atm. The scale reading decreases. B. Two cylinders contain liquids. The pressure, P1, at the bottom of cylinder 1 is higher than the pressure, P2, at the bottom of cylinder 2. The ratio, P1/P2, of the absolute pressures is greater than the ratio of the gauge pressures. C. Water contracts when cooled (at least above 4 °C). A beaker of water is cooled from 50 °C to 20 °C. The pressure at the bottom of the beaker increases.

OpenStudy (anonymous):

A. Probably false. That would have to be some scale to measure the loss in weight due to the loss of air in the duck's lungs. B. True P1/P2 > (P1 + A)/(P2 + A) E.G., (2)/(1) > (2+15)/(1+15)=17/16 C. False. Pressure = force/area = weight/area, unchanged by freezing.